Que 1. Sameer Electronics is producing a large range of electrical goods. It has under construction two projects ‘X’ and ‘Y’ each costing Rs. 120 lakhs.
The projects are mutually exclusive and the company is considering the question of selecting one of the two. Cash flows have been worked out for both the projects and details are given below. X has a life of 8 years and Y has a life of 6 years. Both will have zero salvage value at the end of their operational lives. The company is already making profits and its tax rate is 50%. The cost of capital of the company is 15%.
Net cash inflow (Rs. in lakhs)
Year 1 2 3 4 5 6 7 8
Project X 25 35 45 65 65 55 35 15
Project Y 40 60 80 50 30 20 - -
PV @ 15% 0.870 0.756 0.685 0.572 0.497 0.432 0.376 0.327
The company follows straight line method of depreciation. Advise the company regarding the selection of the project.
1
Expert's answer
2020-06-08T11:08:14-0400

X

"NPV=-120+\frac{25}{1.15}+\frac{35}{1.15^2}+\frac{45}{1.15^3}+\frac{65}{1.15^4}+\frac{65}{1.15^5}+\frac{55}{1.15^6}+\frac{35}{1.15^7}+\frac{15}{1.15^8}=-120+21.74+26.47+29.59+37.16+32.32+23.78+13.16+4.90=-120+189.12=69.12"


Y

"NPV=-120+\frac{40}{1.15}+\frac{60}{1.15^2}+\frac{80}{1.15^3}+\frac{50}{1.15^4}+\frac{30}{1.15^5}+\frac{20}{1.15^6}=-120+34.78+45.37+52.60+28.59+14.92+8.65=64.91"


in 8 years X project is more profitable

If we bring to a full period of 6 years, then the NPV of project X will be 51.06 and then the project will be more profitable